There is a daily demand of 10 units with standard deviation 3 unit.
The lead time is 14 days with the review period is 30 days.
The company set policy of 98% demand satisfaction from items in stock.
Beginning inventory is 150 units.
How many units should be ordered?
Answer #1

Given:

Daily demand = 10 units

Mean monthly demand (d) = 10 *30 units / month = 300 units per month

std deviation (sd) = 3 units per day = 3*30 units / month = 90 units/ month

Lead Time (L) = 14 days

Service level = 98%

z value corresponding to service level 98% = normsinv (0.98) = 2.054

Safety Stock = z value * standard deviation along lead time

Safety Stock (SS) = z* sd = 2.054 * 90 = 184.86 units

Reorder point ( R) = dL + SS = (300*14) + 184.86 = 4200 + 184.86 = 4384.86 units

Beginning inventory (BI) = 150 units

Hence no of units which must be ordered = R – BI = 4384.86 – 150 = 4234.86 units